Assume the following exchange rates: $2 = £1 in New York ¥410 = £1 in London ¥200 = $1 in Tokyo Indicate how profitable triangular, or three-point, arbitrage can take place.


1
Expert's answer
2021-05-03T10:56:53-0400

Lets estimate the rates:"\$ / \pounds = 2, \pounds / \yen = \frac{1}{420}, \yen /\$ =200"


Lets Multiply the rates: "2 \times \frac {1}{420} \times 200 = \frac {400}{420} \cancel {=} 1" therefore there is an arbitrage possibility

1)Lets suppose we have 100 ¥ and we exchange it for $ then "100 \yen = \frac {100}{200} =0.5\$"

2) then exchange 0.5$ for pounds: "0.5\$ = 0.25 \pounds"

3) then exchange it for yen "0.25 \pounds = 0.25 \times 410 = 102,5 \yen"

that means that we earned 2.5 yen from 100 yen. In percentage we gained 2.5%